### v1

# FedSRD: Sparsify-Reconstruct-Decompose for Communication-Efficient Federated Large Language Models Fine-Tuning

Abstract.
The current paradigm of training large language models (LLMs) on publicly available Web data is becoming unsustainable, with high-quality data sources in specialized domains nearing exhaustion. Federated Learning (FL) emerges as a practical solution for the next generation of AI on a decentralized Web, enabling privacy-preserving collaborative fine-tuning by leveraging private data distributed across a global client base. While Low-Rank Adaptation (LoRA) is the standard for efficient fine-tuning, its application in federated settings presents a critical challenge: communication overhead remains a significant bottleneck across the Web’s heterogeneous network conditions. The structural redundancy within LoRA parameters not only incurs a heavy communication burden but also introduces conflicts when aggregating client updates. To address this, we propose FedSRD, a Sparsify-Reconstruct-Decompose framework designed for communication-efficient FL. We first introduce an importance-aware sparsification method that preserves the structural integrity of LoRA updates to reduce the uploaded parameter count. The server then reconstructs and aggregates these updates in a full-rank space to mitigate conflicts. Finally, it decomposes the global update into a sparse low-rank format for broadcast, ensuring a symmetrically efficient cycle. We also propose an efficient variant, FedSRD-e, to reduce computational overhead. Experimental results on 10 benchmarks demonstrate that our framework significantly reduces communication costs by up to 90% while even improving model performance on heterogeneous client data.
Federated Learning, Large Language Models, Communication, LoRA, Fine-tuning
copyright: nonedoi: XXXXXXX.XXXXXXXccs: Computing methodologies Cooperation and coordination
1. Introduction
Large Language Models (LLMs), pre-trained on vast web-scale corpora, have demonstrated remarkable generalist capabilities (Ouyang et al., 2022; Kaplan et al., 2020). However, realizing their full potential for specialized downstream applications requires fine-tuning on diverse, high-quality, domain-specific data. Such data is often proprietary and distributed across numerous private sources, resulting in fragmented ”data silos.” As centralizing this sensitive data is frequently impractical due to privacy regulations and cost constraints, Federated Learning (FL) offers a compelling solution, enabling multiple clients to collaboratively fine-tune a global shared LLM without direct data exchange. This paradigm is central to building the next generation of intelligent, decentralized Web applications.
While FL provides a privacy-preserving solution, fine-tuning billion-parameter LLMs across distributed clients remains computationally and communicatively prohibitive for individual participants. To overcome this, Parameter-Efficient Fine-Tuning (PEFT) methods, particularly Low-Rank Adaptation (LoRA) (Hu et al., 2022), have become the standard (Ding et al., 2023; Kuang et al., 2024; Ye et al., 2024). As illustrated in Figure 1, LoRA significantly reduces the training burden by freezing the pre-trained model weights and optimizing only a small set of trainable, low-rank matrices. In a federated setting, clients train and communicate only these LoRA parameters, achieving performance comparable to full-parameter fine-tuning at a fraction of the resource cost.
Despite the efficiency of LoRA, the iterative communication of its parameters remains a substantial overhead. This becomes a critical bottleneck when deployed across the highly heterogeneous and unpredictable network conditions characteristic of the global Web and mobile systems (Dorfman et al., 2023; Mundt, 2004). High communication costs can exclude clients with limited bandwidth, such as those in rural areas or on mobile devices, undermining data diversity and posing fairness and inclusivity challenges for the global Web ecosystem. While sparsification is a common strategy for reducing communication, conventional methods are ill-adapted for federated LoRA fine-tuning on non-IID data. As shown in Figure 2, naïve approaches like magnitude-based pruning (Aji and Heafield, 2017; Yadav et al., 2023; Deep et al., 2024) disrupt the critical structural relationship between the LoRA matrices $A$ and $B$, while random sparsification (Yu et al., 2024) also leads to performance degradation at high sparsity ratios. Crucially, both methods fail to preserve client-specific knowledge and exacerbate conflicts during server-side aggregation under data heterogeneity. This results in an undesirable trade-off: reduced communication at the expense of significant performance loss. This leads us to our central research question:
How can we significantly reduce communication costs in federated LLMs fine-tuning without sacrificing model performance under heterogeneous client data?
Figure 1. An illustration of Federated LoRA Fine-tuning. Clients train and communicate only the LoRA matrices. However, this payload can still amount to hundreds of megabytes per round, posing a significant bottleneck.
Figure 2. Performance comparison under different sparsity ratios. Both random and magnitude-based sparsification suffer performance decline at high sparsity ratios, motivating the need for a more advanced, structure-aware strategy.
To address this question, we introduce FedSRD, a novel Sparsify-Reconstruct-Decompose framework for communication-efficient federated LLM fine-tuning. Our framework is designed to maintain high model performance through three sequential operations.
First, on the client side, we develop an importance-aware Sparsification strategy. Instead of relying on parameter magnitude, we calculate the importance scores that quantify the contribution of each element in the update matrices ($\Delta B$ and $\Delta A$) to the overall low-rank update ($\Delta W$). This effectively preserves the critical structural interplay between the LoRA matrices. Furthermore, recognizing that different layers in an LLM serve distinct functions, FedSRD adaptively determines the sparsity ratio for each layer based on the kurtosis of its importance score distribution. Aggressive pruning is applied where high importance is concentrated in a few parameters, while conservative pruning is adopted where importance is more diffuse. Second, the server executes the Reconstruction and Decomposition. Upon receiving sparse updates, the server first reconstructs each client’s full-rank weight update ($\Delta W$) before aggregation. This enables aggregation to occur in the full-rank parameter space, a process that is mathematically sound and effectively mitigates the update conflicts arising from non-IID data (Wang et al., 2024; Sun et al., 2024). Finally, to create a symmetrically efficient communication cycle, the server performs a novel Decomposition of the aggregated global update back into a sparse, low-rank format for clients to download. Using a Taylor expansion-based approximation, we employ an alternating scheme that solves for and transmits only one sparse low-rank matrix per round. Furthermore, for scenarios where server-side computation is the bottleneck (e.g., on edge servers), we propose an efficient variant, FedSRD-e, which bypasses the computationally intensive SVD projection during decomposition.
We demonstrate the effectiveness of our framework through extensive experiments that simulate practical non-IID scenarios by finetuning an LLM with clients from diverse specialist domains. Our evaluation on 8 in-domain and 2 out-of-domain benchmarks shows that FedSRD reduces both upload and download communication by 90% while achieving even better performance. In summary, our contributions are:
  • We propose a novel, structure-aware sparsification method for federated LoRA fine-tuning based on importance scores. It effectively identifies and preserves the most salient parameters, overcoming the limitations of conventional sparsification methods.
  • We propose a two-step server-side process: (1) reconstructing and aggregating client updates in a full-rank space to ensure mathematically sound aggregation and conflict mitigation, and (2) an alternating decomposition method to create sparse global updates for efficient download. We also introduce an efficient variant of this process to reduce server computation.
  • We conduct extensive experiments demonstrating that our methods can save over 90% of communication costs while achieving superior model performance, validating FedSRD as a practical and high-performing solution for federated LLMs fine-tuning on the Web.
2. Related Works
2.1. Parameter-Efficient Fine-Tuning
The large scale of modern LLMs makes the full-parameter fine-tuning computationally prohibitive in many resource-constrained environments, such as on mobile or edge devices. To address this, Parameter-Efficient Fine-Tuning (PEFT) methods (Ding et al., 2023) have been developed, which freeze most of a pre-trained model’s parameters and only update a small part. Current approaches include adapter-based methods (He et al., 2021; Edalati et al., 2025; Houlsby et al., 2019; Hu et al., 2022) that insert small, trainable modules between model layers, and prompting methods (Petrov et al., 2023; Lester et al., 2021), such as prefix-tuning (Li and Liang, 2021), that add learnable tokens to the input sequence. Among these, Low-Rank Adaptation (LoRA) (Hu et al., 2022) has emerged as a dominant and effective approach. The key insight of LoRA is that weight updates during fine-tuning possess a low intrinsic rank. LoRA represents the update for each target weight matrix as the product of two smaller, trainable low-rank matrices, significantly reducing the number of trainable parameters and associated memory footprint while achieving comparable performance.
2.2. Federated LLMs Fine-tuning
Federated Learning (FL) enables the collaborative fine-tuning of LLMs on decentralized private data, addressing both data scarcity and privacy concerns. A common approach is to combine PEFT methods with standard aggregation algorithms (Long et al., 2024; Kuang et al., 2024; Zhang et al., 2024; Nguyen et al., 2024; Ye et al., 2024). Recent studies point out that this naive combination introduces aggregation inaccuracy, where separately averaging the decomposed LoRA matrices ($B$ and $A$) from clients is not mathematically equivalent to averaging the true underlying model updates ($BA$). Several advanced methods have been proposed to mitigate this issue. FFA-LoRA (Sun et al., 2024) ensures a mathematically sound aggregation by freezing the LoRA $A$ matrices and only aggregating the $B$ matrices. FLoRA (Wang et al., 2024) stacks clients’ LoRA matrices to ensure accurate aggregation. FlexLoRA (Bai et al., 2024) proposes to aggregate the true model updates ($BA$) on the server and then employs Singular Value Decomposition (SVD) to derive new global low-rank matrices for clients. FedICU (Liao et al., [n. d.]) disentangles client updates into consensus and divergence components, which are then aggregated separately.
2.3. Communication-Efficient Federated LLMs Finetuning
Communication overhead remains a primary bottleneck in FL systems (Yan et al., 2025, 2024; Sattler et al., 2019), a problem that is exacerbated when fine-tuning LLMs over the public internet. While PEFT methods provide a baseline level of efficiency, further optimizations are crucial. Structured methods impose constraints on the LoRA updates. For instance, FFA-LoRA (Sun et al., 2024) halves the communication cost by transmitting only the $B$ matrices. Building on this, FedSVD (Lee et al., 2025) iteratively re-initializes LoRA $A$ matrices. Fed-SB (Singhal et al., 2025) freezes the LoRA matrices entirely and instead inserts a small, trainable square matrix for aggregation. While these approaches reduce communication, they often do so at the cost of model performance by constraining the expressive capacity of LoRA. Other lines of work focus on generic compression. Quantization methods (Aji and Heafield, 2017; Konečnỳ et al., 2016) represent parameters with fewer bits, but they often fail to achieve high compression ratios and can lead to significant accuracy degradation, particularly in non-IID settings. Sparsification methods, which transmit only a subset of non-zero parameters, can achieve much higher compression ratios (Aji and Heafield, 2017; Sahu et al., 2021). Yet, traditional approaches are not directly applicable to LoRA. Magnitude-based pruning (Aji and Heafield, 2017) is suboptimal because it ignores the coupled relationship between the two LoRA matrices. Similarly, random sparsification (Yang et al., 2025; Yu et al., 2024) fails to preserve critical parameters, resulting in inferior performance at high sparsity ratios.
3. Preliminaries
3.1. LoRA in LLMs Fine-tuning
LoRA has been proven to exhibit better performance than other PEFT methods in the federated learning system (Kuang et al., 2024). LoRA constrains the weight update in low-rank space:
(1) $W_{0}+\Delta W=W_{0}+BA,$
where $W_{0}$ is the frozen model weights, $B\in\mathbb{R}^{d_{out}\times r}$ and $A\in\mathbb{R}^{r\times d_{in}}$ are paired low-rank trainable matrices inserted in the linear weights to approximate the update $\Delta W\in\mathbb{R}^{d_{out}\times d_{in}}$, where $r\ll\min(d_{out},d_{in})$. $B$ is initialized to zero and $A$ is initialized with random Gaussian noises (Hu et al., 2022).
3.2. Problem Formulation
We consider an FL system with one server and $m$ clients. Each client has its local private dataset $\mathcal{D}^{i}=\{x_{j}^{i},y_{j}^{i}\}_{j=1}^{|\mathcal{D}^{i}|}$. These datasets may originate from different domains or tasks, creating a heterogeneous (non-IID) scenario. All clients collaboratively fine-tune a shared LLM $\mathcal{M}$ to perform well on their respective data distributions. This can be formulated as the following optimization problem:
(2) $\min_{\mathcal{M}}\mathcal{L}(\mathcal{M})=\sum_{i}^{m}\mathcal{L}_{i}(\mathcal{M},\mathcal{D}^{i}),$
where $\mathcal{L}_{i}(\mathcal{M},\mathcal{D}^{i})$ represents the loss of the LLM $\mathcal{M}$ on local dataset $\mathcal{D}^{i}$. Following the recent studies, we adopt LoRA to fine-tune the LLM $\mathcal{M}$ in local fine-tuning on each client. The local optimization of client $i$ could be formulated as:
(3) $\min\mathcal{L}_{i}=\min_{\mathcal{P}(\mathcal{M})}\mathbb{E}_{(x_{j}^{i},y_{j}^{i})\sim\mathcal{D}^{i}}L(x_{j}^{i},y_{j}^{i},\mathcal{P}(\mathcal{M})),$
where $\mathcal{P}(\mathcal{M})$ denotes the trainable parameters set of LoRA. $L$ is the cross-entroy loss on completion $y_{j}^{i}$ given $x_{j}^{i}$.
4. Proposed Method
In this section, we introduce FedSRD, a Sparsify-Reconstruct-Decompose framework for federated LoRA fine-tuning, illustrated in Figure 3. FedSRD establishes a symmetrically efficient communication cycle by transmitting only sparse updates of the LoRA parameters. At each communication round $t$, a client $i$ first computes its local LoRA updates $\Delta B_{i}^{t}$ and $\Delta A_{i}^{t}$, then uses our proposed importance-aware strategy to sparsify them for upload (Section 4.1). The server then reconstructs each client’s full weight matrix and aggregates them in the full-rank space to mitigate non-IID conflicts (Section 4.2). Finally, to ensure an efficient download, the server decomposes the resulting global update into a single sparse low-rank matrix for broadcast using an alternating, Taylor approximation-based method (Section 4.3). The pseudo-code is provided in Appendix A.
Figure 3. The FedSRD framework. (1) Client-Side sparsification: Each client computes its LoRA updates $\Delta B_{i}^{t}$ and $\Delta A_{i}^{t}$ and uses our importance-aware sparsification to generate sparse updates $\Delta B_{i,s}^{t}$ and $\Delta A_{i,s}^{t}$ for upload. (2) Server-Side reconstruction to aggregation: The server reconstructs each client’s full weight matrix $W_{i}^{t}$ and aggregates them in the full-rank space. (3) Server-Side decomposition: The server computes the global update $\Delta W^{t}$, which is then decomposed into a single sparse matrix for efficient download.
4.1. Importance-aware Sparsification
Motivation
To reduce upload communication costs, sparsification is a natural approach. However, mainstream methods are ill-suited for federated LoRA fine-tuning. Magnitude-based sparsification is ineffective as it prunes each LoRA matrix independently, ignoring the crucial structural interplay between $B$ and $A$. On non-IID data, random-based sparsification fails to preserve performance at high sparsity ratios because it cannot differentiate between critical and unimportant parameters. This motivates a sparsification method that can (1) account for the structural relationship between $B$ and $A$ and (2) robustly preserve the most salient parameters.
To address these challenges, we propose importance-aware sparsification. Instead of transmitting the full LoRA matrices, we transmit only their updates at each round. Specifically, on client $i$ at round $t$, we first conduct LoRA fine-tuning on its local dataset to obtain the new LoRA matrices $B_{i}^{t}$ and $A_{i}^{t}$, where
(4) $B_{i}^{t}=B_{i}^{t-1}+\Delta B_{i}^{t},\quad A_{i}^{t}=A_{i}^{t-1}+\Delta A_{i}^{t},$
where $B_{i}^{t-1}$ and $A_{i}^{t-1}$ are the matrices from the previous round  [^1] Since our operations are layer independent, we omit layer-specific notations for clarity.. The full-rank weight update $\Delta W_{i}^{t}$ can be expressed as:
(5) $\Delta W_{i}^{t}$ $=B_{i}^{t}A_{i}^{t}-B_{i}^{t-1}A_{i}^{t-1}$
$=B_{i}^{t}(A_{i}^{t-1}+\Delta A_{i}^{t})-B_{i}^{t-1}A_{i}^{t-1}$
$=(B_{i}^{t}-B_{i}^{t-1})A_{i}^{t-1}+B_{i}^{t}\Delta A_{i}^{t-1}$
$=\Delta B_{i}^{t}A_{i}^{t-1}+B_{i}^{t}\Delta A_{i}^{t}.$
This formulation reveals that the true contribution of an element in $\Delta B_{i}^{t}$ depends on the corresponding row in $A_{i}^{t-1}$, and likewise for $\Delta A_{i}^{t}$ and $B_{i}^{t}$. To formalize this, we therefore define importance score $I(\cdot)$ to measure an element’s contribution to the final weight update by scaling its magnitude with the norm of the corresponding vector it interacts with in the other matrix:
(6) $I(\Delta B_{i}^{t}[u,v])$ $=|\Delta B_{i}^{t}[u,v]|\cdot\|A_{i}^{t-1}[v,:]^{T}\|_{2},$
$I(\Delta A_{i}^{t}[u,v])$ $=|\Delta A_{i}^{t}[u,v]|\cdot\|B_{i}^{t}[:,u]||_{2}.$
A higher importance score indicates a greater contribution to the final weight update. We explain the rationale behind in Appendix B. We then prune elements with the lowest scores, transmitting only the most important portion of the parameters.
Furthermore, recognizing that LoRA matrices in different layers play distinct roles (Luo et al., 2025), we propose to adaptively assign a sparsity ratio $\rho$ to each matrix. The core intuition is to use the kurtosis of the importance score distribution reveals a layer’s functional ”prunability”. A high kurtosis (a leptokurtic, heavy-tailed distribution) signifies that importance is concentrated in a few parameters, justifying an aggressive sparsity ratio. Conversely, a low kurtosis (a platykurtic, flat distribution) implies that importance is diffused, necessitating a more conservative ratio. The layer-specific sparsity ratio is thus calculated as:
[…]
where $\alpha$ is a hyperparameter. The sparsity ratios are truncated if they exceed a pre-set upper bound to ensure validity. We provide a detailed explanation and justification of this design in Appendix F. We use the adaptive sparsity ratios to prune $\Delta B_{i}^{t}$ and $\Delta A_{i}^{t}$:
[…]
where the $\Delta B_{i,s}^{t}$ and $\Delta A_{i,s}^{t}$ are sparsified matrices for uploading. The client can upload the sparse matrices in a compressed format.
4.2. Reconstruct to Aggregation
Motivation
Most existing federated learning methods like FedAvg (McMahan et al., 2017), when applied to LoRA, average the clients’ LoRA matrices independently. This aggregation is not mathematically sound and introduces unexpected noises because it multiplicative relationship between the matrices and fails to mitigate update conflicts arising from non-IID client data (Sun et al., 2024; Bai et al., 2024; Wang et al., 2024). This motivates us to develop an advanced aggregation method.
To enable a mathematically sound aggregation that preserves client knowledge, we propose to first reconstruct each client’s update in the full-rank space. At round $t$, upon receiving the sparse updates $\Delta B_{i,s}^{t}$ and $\Delta A_{i,s}^{t}$ from $m$ participating clients, the server first computes each client’s new LoRA matrices, $B_{i}^{t}$ and $A_{i}^{t}$, using Equation 4. It then reconstructs each client’s full-rank weight matrix and aggregates them:
(9) $W^{t}=\frac{1}{m}\sum_{i}^{m}B_{i}^{t}A_{i}^{t},$
where $W^{t}$ is the aggregated LoRA weights in the full-rank parameter space. By performing aggregation in this higher-dimensional space, the client-specific knowledge encoded in $B_{i}^{t}$ and $A_{i}^{t}$ is properly combined, reducing the likelihood of destructive interference.
4.3. Alternative Decomposition
Motivation
After aggregating in the full-rank space, the server must decompose the result back into low-rank matrices for the clients. Previous works (Bai et al., 2024; Lee et al., 2025) often use Singular Value Decomposition (SVD) to construct new dense matrices $B^{t}$ and $A^{t}$. However, transmitting these dense matrices would create a severe download bottleneck, negating the benefit of a sparse upload. Naively sparsifying the full state matrices $B^{t}$ and $A^{t}$ would damage the learned model, and one cannot simply compute an update as $B^{t}-B^{t-1}$ because it does not correctly approximate the global update ($B^{t}A^{t}-B^{t-1}A^{t-1}\neq(B^{t}-B^{t-1})A^{t-1}$). This motivates a novel decomposition method that can produce a sparse update for download.
Our goal is to find sparse updates $\Delta B^{t}$ and $\Delta A^{t}$ that approximate the change from the previous global state $W^{t-1}$ to the new one $W^{t}$. First, we project $W^{t}$ to its $r$-rank approximation via SVD:
(10) $U,S,V_{h}$ $=SVD(W^{t}),$
$W^{t}_{low\_rank}$ $=U[:,:r]S[:r,:r]V_{h}[:r,:],$
where $W^{t}_{low\_rank}$ is the $r$-rank approximation of $W^{t}$, $r$ is the LoRA rank. This process inherently denoises the aggregated result by discarding less important components and retaining the consensus. Assuming the global states are low-rank, we can state that $\exists B^{t-1},A^{t-1}$ such that $W^{t-1}_{\text{low\_rank}}=B^{t-1}A^{t-1}$, and we seek $\Delta B^{t},\Delta A^{t}$ such that $W^{t}_{\text{low\_rank}}=(B^{t-1}+\Delta B^{t})(A^{t-1}+\Delta A^{t})$. We can approximate the update using a first-order Taylor expansion:
(11) $W$ ${}^{t}_{low\_rank}-W^{t-1}_{low\_rank}$
$=(B^{t-1}+\Delta B^{t})(A^{t-1}+\Delta A^{t})-B^{t-1}A^{t-1}$
$=B^{t-1}\Delta A^{t}+\Delta B^{t}A^{t-1}+\Delta B^{t}\Delta A^{t}$
$\approx B^{t-1}\Delta A^{t}+\Delta B^{t}A^{t-1},$
where the higher-order term $\Delta B^{t}\Delta A^{t}$ is omitted (discussed in Appendix G). To solve this underdetermined system, we alternate between setting one of the deltas to zero. In even-numbered rounds, we set $\Delta B^{t}=\mathbf{0}$ and solve for $\Delta A^{t}$:
(12) $W^{t}_{low\_rank}-W^{t-1}_{low\_rank}=B^{t-1}\Delta A^{t},$
$\Delta A^{t}=(B^{t-1})^{\dagger}(W^{t}_{low\_rank}-W^{t-1}_{low\_rank}),$
where $(B^{t-1})^{\dagger}$ is the Moore-Penrose pseudoinverse. In odd-numbered rounds, we set $\Delta A^{t}=\mathbf{0}$ and solve for $\Delta B^{t}$:
(13) $W^{t}_{low\_rank}-W^{t-1}_{low\_rank}=\Delta B^{t}A^{t-1},$
$\Delta B^{t}=(W^{t}_{low\_rank}-W^{t-1}_{low\_rank})(A^{t-1})^{\dagger},$
The resulting single update matrix ($\Delta A^{t}$ or $\Delta B^{t}$) can be simply sparsified using random pruning and rescaled (Yu et al., 2024) to produce a sparse update ($\Delta A^{t}_{s}$ or $\Delta B_{s}^{t}$) for download. Clients receive this single sparse matrix and update their local model accordingly, completing the communication-efficient cycle.
FedSRD-e: a Computationally Efficient Variant
For scenarios where server-side computation is the primary bottleneck, such as in edge devices, we propose a computationally efficient variant. This variant, FedSRD-e, bypasses the SVD projection and instead directly approximates the full-rank global update, $\Delta W^{t}=W^{t}-W^{t-1}$. This matrix $\Delta W^{t}$ replaces $W^{t}_{\text{low\_rank}}-W^{t-1}_{\text{low\_rank}}$ on the left-hand side of Equation 11. The subsequent steps of solving for a single update matrix remain the same. This variant significantly reduces computation by avoiding the costly SVD operation. It only consumes slightly more time than the simple FedAvg method, serving as a powerful, high-speed alternative.
Remark
Note that both the reconstruction and decomposition operations are layer-independent, meaning they can be performed layer by layer. Thus, the additional memory resource demand on the server is negligible.
Table 1. Details of the local datasets used by each client. All datasets are from the training split.
Domain Dataset Task Type
Code Code Alpaca Code Generation
Medical MedQA Medical Question Answering
Finance Finance Alpaca Financial Question Answering
Math GSM8K (train) Mathematical Reasoning &
Math (train) Math Problems Solving
Table 2. Details of the benchmarks used in our evaluation.
Category Domain Benchmark # Instance $k$-Shot
In-domain Code HumanEval (Chen et al., 2021) 164 0
Sanitized MBPP (Austin et al., 2021) 257 3
Medical MedQA (Jin et al., 2021) 1,273 1
MedMCQA (Pal et al., 2022) 4,183 1
Finance FinEval (Zhang et al., 2023) 4,661 (34 sub-tasks) 0
FinanceIQ (Zhang and Yang, 2023) 7,173 (10 sub-tasks) 0
Math GSM8K (Cobbe et al., 2021) 1319 0
MATH (Hendrycks et al., 2021) 5000 0
Out-of-domain General AGIEval (Zhong et al., 2023) 8,062 (20 sub-tasks) 0
Law LawBench (Fei et al., 2023) 10,000 (20 sub-tasks) 1
5. Experiments
5.1. Experimental Setup
5.1.1. Datasets
We simulate an FL system with 4 clients from distinct specialist domains: Code, Medical, Finance, and Math. We use CodeAlpaca (Chaudhary, 2023), the training set of MedQA (Jin et al., 2021), Finance Alpaca [^2] https://huggingface.co/datasets/gbharti/finance-alpaca, and a combination of the training sets from GMS8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) as the local datasets for these clients, respectively.
5.1.2. Baselines
We compare our framework FedSRD and FedSRD-e with 6 state-of-the-art federated learning methods: (1) FedAvg (McMahan et al., 2017), (2) FedProx (Li et al., 2020), (3) FFA-LoRA (Sun et al., 2024), which freezes LoRA $A$ matrcies to ensure efficient communication and effective aggregation, (4) FlexLoRA (Bai et al., 2024), which aggregates in full-rank space and uses SVD for download, (5) FedICU (Liao et al., [n. d.]), which selectively updates on global model and perform consensus-divergence splitting in the server-side aggregatio, (6) Fed-DARE, which applies random sparsification with rescaling (DARE (Yu et al., 2024)) to LoRA updates for upload. A detailed description of baselines is in Appendix C.
5.1.3. Benchmarks
We evaluate baselines and our methods on 10 benchmarks, including 8 in-domain benchmarks from Code, Medical, Finance, and Math domains, and 2 out-of-domain benchmarks to evaluate the knowledge retention of other domains during fine-tuning. We provide details of these benchmarks in Table 2.
5.1.4. Implementation
In our experiments, we adopt Llama3.2-3B  [^3] https://huggingface.co/meta-llama/Llama-3.2-3B and Qwen2-7B (Team, 2024) as base models for federated fine-tuning over 50 rounds. We apply LoRA to all linear layers with rank $r=64$ for Llama3.2-3B and $r=32$ for Qwen2-7B. The learning rate is tuned from 1e-5 to 2e-4. For Fed-DARE, the sparsity ratio is 0.9 as suggested (Yu et al., 2024). For our method, we tune $\alpha$ in Equation 7 within [0.85, 0.9] and set the download sparsity ratio to 0.8. For other hyperparameters, we follow the suggested values from the original papers. We report average performance on in-domain and out-of-domain benchmarks, as well as the per-round communication cost per client. We conduct training on NVIDIA A800 GPUs and inference on NVIDIA GeForce RTX 3090 GPUs with bfloat16 precision.
Table 3. Comparison of model performance and per-round communication cost using Llama3.2-3B as the base model. OOD represents Out-of-Domain. The best and second-best results are highlighted in bold and underline, respectively.
Method
Comm.
Cost
Average In-Domain Out-of-Domain
In-Domain OOD Code Medical Finance Math AGIEval Law
HumanEval MBPP MedQA MedMCQA FinEval FinanceIQ GSM8K MATH
FedAvg 742 33.90 24.36 26.85 47.86 40.45 41.53 41.96 38.08 28.96 5.56 22.47 26.24
FedProx 742 33.43 23.82 28.05 48.25 35.98 40.06 42.75 37.82 28.05 5.70 21.58 26.06
FFA-LoRA 371 28.39 26.12 3.05 46.30 39.95 40.11 40.23 36.02 24.79 5.10 24.63 27.60
FlexLoRA 742 32.36 26.08 28.75 47.47 36.29 41.24 42.22 37.09 27.37 5.68 24.92 27.24
FedICU 568 33.91 25.76 31.10 49.42 40.15 41.07 40.83 37.56 25.09 5.26 24.77 26.75
Fed-DARE 190 33.57 24.25 26.60 48.25 39.28 40.04 40.40 38.04 26.05 5.26 21.94 26.56
FedSRD-e 74 34.26 27.95 29.27 45.87 41.63 41.17 42.22 39.47 30.48 5.88 28.81 27.08
FedSRD 74 35.12 26.17 31.10 49.42 41.95 41.62 43.44 39.13 28.96 5.36 25.54 26.79
Table 4. Comparison of model performance and per-round communication cost using Qwen2-7B as the base model. OOD represents Out-of-Domain. The best and second-best results are highlighted in bold and underline, respectively.
Method
Comm.
Cost
Average In-Domain Out-of-Domain
In-Domain OOD Code Medical Finance Math AGIEval Law
HumanEval MBPP MedQA MedMCQA FinEval FinanceIQ GSM8K MATH
FedAvg 616 57.04 41.02 59.15 37.35 61.67 75.23 65.16 49.04 64.29 44.46 48.25 33.79
FedProx 616 58.27 40.72 63.42 36.58 62.13 77.34 65.77 49.96 66.26 44.72 48.62 32.82
FFA-LoRA 308 53.36 43.41 68.29 17.90 51.55 57.02 67.76 48.92 69.37 46.08 46.92 39.89
FlexLoRA 616 55.89 40.49 63.41 35.02 58.83 68.30 67.68 50.53 58.61 44.70 47.81 33.17
FedICU 472 57.98 42.56 60.37 38.91 60.33 78.72 65.86 48.83 66.19 44.62 48.44 36.28
Fed-DARE 158 57.01 41.72 60.37 31.91 61.12 75.30 66.03 50.84 65.96 44.58 48.37 35.07
FedSRD-e 61 61.13 42.59 68.09 36.56 77.80 78.10 64.21 48.90 70.20 45.20 45.82 39.35
FedSRD 61 61.19 42.58 68.29 36.19 77.77 78.94 64.55 48.48 69.93 45.36 45.93 39.22
Figure 4. Average in-domain performance vs. per-round communication cost (Llama3.2-3B). Our methods are Pareto optimal, achieving the highest performance with the lowest communication cost.
Figure 5. Average in-domain performance vs. per-round communication cost (Qwen2-7B). Our methods are Pareto optimal, achieving the highest performance with the lowest communication cost.
5.2. Experimental Results
5.2.1. Performance Comparison
The main results, reported in Table 3 (Llama3.2-3B) and Table 4 (Qwen2-7B), comprehensively demonstrate the effectiveness and efficiency of our FedSRD framework. We analyze performance from three key perspectives.
In-Domain Performance.
Across the experiments, our proposed methods, FedSRD and FedSRD-e, consistently achieve state-of-the-art performance on in-domain benchmarks. As shown in Table 3, with the Llama3.2-3B model, FedSRD and FedSRD-e achieve the highest average in-domain performance score, outperforming all baseline methods. This trend is further confirmed with the larger Qwen2-7B model in Table 4, where FedSRD and FedSRD-e again lead with top performance, respectively. While some baselines may achieve peak performance on specific benchmarks, they often falter in other domains. In contrast, our framework exhibits robust and consistently high performance across all four specialized domains, underscoring its superior capability in aggregating diverse knowledge from specialist clients to build powerful global LLMs.
Communication Efficiency.
A standout advantage of our framework is its remarkable communication efficiency. FedSRD and FedSRD-e reduce the per-round communication overhead to just 74 MB for Llama3.2-3B and 61 MB for Qwen2-7B. This represents a reduction of $\approx$ 90% compared to standard methods like FedAvg (742 MB) and advanced methods like FlexLoRA (742 MB). This reduction is achieved by our proposed importance-aware sparsification for uploads and alternative decomposition for downloads (see Appendix D for a detailed analysis). Crucially, this significant gain in efficiency does not come at the cost of performance. On the contrary, FedSRD and FedSRD-e achieve higher performance than baselines, directly answering our central research question. This is visualized in Figures 4 and 5, where our methods are shown to be Pareto optimal: achieving the best performance with the least communication.
Out-of-Domain Performance.
Our methods also demonstrate strong performance on out-of-domain (OOD) benchmarks, indicating effective mitigation of catastrophic forgetting. For Llama3.2-3B, FedSRD and FedSRD-e achieve the leading OOD performance. For Qwen2-7B, they also maintain competitive performance. Although FFA-LoRA also exhibits strong OOD retention, we attribute this to its mechanism of freezing the LoRA $A$ matrices. which inherently constrains the capacity and hinders new knowledge acquisition in favor of preserving prior information. In contrast, our methods provide a more balanced and consistently high level of performance across both in-domain and out-of-domain evaluations. Further, we posit that the SVD projection in FedSRD, by seeking a low-rank consensus for in-domain tasks, may inadvertently filter out some general knowledge from the base model, which could account for the superior OOD performance of FedSRD-e. In summary, our framework not only excels at integrating specialized knowledge but also effectively preserves the general capabilities of the base model.
5.2.2. Abation Study
To validate each component of our FedSRD framework, we conducted an ablation study on Llama3.2-3B, with results in Table 5. Starting with the full model, we sequentially removed or replaced key components. Disabling the adaptive sparsity ratio (using a fixed ratio instead) causes a slight performance drop, affirming the benefit of dynamic adjustment. Replacing our importance-aware sparsification with a random baseline (Yu et al., 2024) leads to a significant degradation, highlighting that our decomposition method effectively preserves aggregated knowledge while reducing download communication costs. Further, ablating our alternative decomposition (reverting to dense SVD download) also harms performance, confirming its effectiveness in preserving knowledge while reducing download costs. Finally, removing the reconstruction step (reverting to naive LoRA matrix averaging) results in severe performance loss, proving that full-rank aggregation is crucial for mitigating client heterogeneity. These results confirm that each component plays an indispensable and synergistic role in achieving the dual objectives of high model performance and extreme communication efficiency. Additionally, we show in Appendix E that our importance-aware sparsification also works as a more effective plugin to enhance conventional methods like FedAvg.
Table 5. Ablation study on Llama3.2-3B. Components are removed sequentially. Note that importance-aware sparsification is substituted with the random sparsification baseline (Yu et al., 2024) for its ablation.
Method In-Domain OOD
FedSRD 35.12 26.17
- adaptive sparsity ratio (Eq. 7) 35.05 25.94
   - importance-aware sparsification (Sec. 4.1) 32.93 24.76
    - alternative decomposition (Sec. 4.3) 32.48 25.73
     - reconstruct (Sec. 4.2) 32.32 23.90
Figure 6. In-domain and out-of-domain (OOD) performance under different base sparsity values ($\alpha$) for FedSRD. Performance is robust across a range of sparsity levels, consistently outperforming the best baseline results.
Table 6. In-domain and out-of-domain (OOD) performance under different download sparsification for FedSRD. Param. Reduction denotes the reduction ratio of the number of parameters downloaded per round.
Download Sparse Ratio In-Domain OOD Param. Reduction
60% 34.32 26.76 80%
80% 35.12 26.17 90%
90% 34.57 25.13 95%
5.2.3. Impact of Sparsification Ratios
We first analyze the uplink base sparsity ratio, $\alpha$, in our importance-aware sparsification. As illustrated in Figure 6, FedSRD is robust and consistently outperforms all baselines even at an extreme sparsity of 98% ($\alpha$=0.95). Interestingly, OOD performance improves with a higher $\alpha$. This suggests that uploading and aggregating fewer parameters helps preserve the pretrained model’s generalizable knowledge by filtering out potentially conflicts in client-specific updates, thus preventing catastrophic forgetting. We can also observe that smaller $\alpha$ values (corresponding to lower sparsity) do not necessitate better performance, since preserving more parameters may introduce irrelevant or noisy knowledge for aggregation. A properly set $\alpha$ therefore offers a dual advantage: it significantly reduces communication costs while simultaneously reducing the potential interference for more effective aggregation.
For the downlink, we apply random sparsification to the decomposed matrices. Table 6 shows that our method could even achieve better performance under aggressive sparsification. These results confirm that FedSRD achieves highly efficient communication and improved performance for federated LLMs fine-tuning.
5.3. Computation Time Analysis
We measured the wall-clock time required for a single FL round in our simulation environment for both the Llama3.2-3B and Qwen2-7B models, with results reported in Table 7. Among the baselines, FedAvg provides a baseline for time consumption. FedProx requires slightly more time due to its proximal term, whereas FFA-LoRA is faster as it only trains and aggregates half of the LoRA matrices. FlexLoRA consumes more time performing SVD for matrix decomposition, and FedICU is the most time-intensive due to its computationally inefficient disentangled aggregation. For our methods, while the full FedSRD requires more time for its $r$-rank SVD projection, our efficient variant, FedSRD-e, introduces negligible additional computation compared to FedAvg, while exhibiting performance second only to the full FedSRD. This result highlights that our FedSRD-e variant achieves a Pareto-optimal balance across model performance, communication efficiency, and time consumption, making it a highly practical and scalable solution for real-world federated LLM fine-tuning.
Table 7. Computation time per round for different methods.
Time per Round (min) Llama3.2-3B Qwen2-7B
FedAvg 1.92 3.06
FedProx 2.10 3.18
FFA-LoRA 1.86 3.00
FlexLoRA 7.20 12.60
FedICU 30.36 51.60
Fed-DARE 1.92 3.06
FedSRD 7.92 13.50
FedSRD-e 2.10 3.15
6. Conclusion
In this paper, we address the critical challenge of communication overhead in federated LoRA fine-tuning for LLMs. We propose FedSRD, a novel Sparsify-Reconstruct-Decompose framework for communication-efficient FL. Its core components include a structure-aware, importance-based sparsification strategy on the client side; a conflict-mitigating reconstruction and aggregation step in the full-rank space on the server side; and a novel alternating decomposition method to produce sparse global updates for efficient download. We also introduce FedSRD-e, a computationally light variant for resource-constrained environments. Through extensive experiments on diverse benchmarks, we demonstrated that FedSRD reduces communication costs by over 90% while simultaneously improving model performance compared to state-of-the-art methods. Our work presents a practical and high-performing solution that significantly lowers the barrier for scalable, collaborative AI development on the decentralized Web.
Appendix A Pseudo-Code of FedSRD and FedSRD-e
Algorithm 1 FedSRD and FedSRD-e Framework
1:Server Initializes: Global LoRA matrices $B^{0},A^{0}$. Let $\mathcal{C}$ be the set of all clients.
2:Initial Synchronization: Each client $i\in\mathcal{C}$ sets its local state $B_{i}^{0}\leftarrow B^{0},A_{i}^{0}\leftarrow A^{0}$.
3:for each communication round $t=1,2,\dots,T$ do
4:  if $t>1$ then
5:   Server broadcasts $\Delta B_{s}^{t-1}$ or $\Delta A_{s}^{t-1}$.
6:   for each client $i\in\mathcal{C}$ in parallel do
7:     Client updates local model: $B_{i}^{t-1}\leftarrow B_{i}^{t-2}+\Delta B_{s}^{t-1}$, $A_{i}^{t-1}\leftarrow A_{i}^{t-2}+\Delta A_{s}^{t-1}$.
8:  for each client $i\in\mathcal{C}$ in parallel do
9:   $\Delta B_{i,s}^{t},\Delta A_{i,s}^{t}\leftarrow\text{ClientTraining}(i)$
10:  Server-Side Reconstruction and Aggregation:
11:  for each client $i\in\mathcal{C}$ do
12:   Client reconstructs its new state $B_{i}^{t},A_{i}^{t}$ using Eq. 4
13:   Server computes client’s full-rank matrix: $W_{i}^{t}\leftarrow B_{i}^{t}A_{i}^{t}$.
14:  Aggregate in full-rank space: $W^{t}\leftarrow\frac{1}{|\mathcal{C}|}\sum_{i}W_{i}^{t}$.
15:  Server-Side Alternating Decomposition:
16:  $W^{t}_{r}\leftarrow\text{SVD}_{r}(W^{t})$. $\triangleright$ FedSRD-e bypasses this SVD step.
17:  Let $\Delta W^{t}_{update}=W^{t}_{r}-B^{t-1}A^{t-1}$. $\triangleright$ For FedSRD-e, use $W^{t}$ instead of $W^{t}_{r}$.
18:  if $t$ is even then
19:   $\Delta A^{t}\leftarrow(B^{t-1})^{\dagger}(\Delta W^{t}_{update})$
20:   $\Delta B^{t}\leftarrow\mathbf{0}$
21:  else
22:   $\Delta B^{t}\leftarrow(\Delta W^{t}_{update})(A^{t-1})^{\dagger}$
23:   $\Delta A^{t}\leftarrow\mathbf{0}$
24:  Update global state on server: $B^{t}\leftarrow B^{t-1}+\Delta B^{t}$, $A^{t}\leftarrow A^{t-1}+\Delta A^{t}$.
25:  Sparsify for download: $\Delta B_{s}^{t},\Delta A_{s}^{t}\leftarrow\text{DARE}(\Delta B^{t},\Delta A^{t})$.
26:
27:function ClientTraining(ii)
28:  Fine-tune model on local data $\mathcal{D}_{i}$ to get new $B^{\prime}_{i},A^{\prime}_{i}$.
29:  Compute local updates: $\Delta B_{i}^{t}\leftarrow B^{\prime}_{i}-B_{i}^{t-1}$, $\Delta A_{i}^{t}\leftarrow A^{\prime}_{i}-A_{i}^{t-1}$.
30:  Importance-Aware Sparsification:
31:  Calculate importance scores $I(\Delta B_{i}^{t}),I(\Delta A_{i}^{t})$ using Eq. 6.
32:  Calculate adaptive sparsity ratios $\rho_{\Delta B_{i}^{t}},\rho_{\Delta A_{i}^{t}}$.
33:  Prune updates to get sparse $\Delta B_{i,s}^{t},\Delta A_{i,s}^{t}$ adaptively.
34:  return $\Delta B_{i,s}^{t},\Delta A_{i,s}^{t}$ to server.
Appendix B Rationale of the Importance Score Calculation
We provide the mathematical motivation for the importance score defined in Equation 6. The core idea is to quantify how much a single parameter’s change in a LoRA update matrix ($\Delta B_{i}^{t}$ or $\Delta A_{i}^{t}$) contributes to the overall change in the full-rank weight matrix ($\Delta W_{i}^{t}$).
Let’s first recall the expression of the weight update:
(14) $\Delta W_{i}^{t}=\Delta B_{i}^{t}A_{i}^{t-1}+B_{i}^{t}\Delta A_{i}^{t}$
We will analyze the contribution of a single element in $\Delta B_{i}^{t}$ to the term $\Delta B_{i}^{t}A_{i}^{t-1}$. The analysis for $\Delta A_{i}^{t}$ is symmetric.
A fundamental property of matrix multiplication is that the product of two matrices, $W=BA$, can be expressed as the sum of the outer products of the columns of $B$ and the rows of $A$. If $B\in\mathbb{R}^{d\times r}$ has columns $\mathbf{b}_{j}$ and $A\in\mathbb{R}^{r\times k}$ has rows $\mathbf{a}_{j}^{T}$, then:
(15) $W=BA=\sum_{j=1}^{r}\mathbf{b}_{j}\otimes\mathbf{a}_{j}^{T}$
This expansion reveals the structural interplay between the two matrices: the final matrix $W$ is constructed from pairs of columns from $B$ and rows from $A$.
Now, let’s consider the contribution of a single scalar element, $(\Delta B_{i}^{t})_{uv}$ (the element in row $u$ and column $v$), to the matrix product $\Delta B_{i}^{t}A_{i}^{t-1}$. We can isolate this contribution by considering a change in only this single element. The effect of this single element on the product can be represented as:
(16) $(\Delta B_{i}^{t})_{uv}\cdot(\mathbf{e}_{u}\otimes\mathbf{e}_{v}^{T})\cdot A_{i}^{t-1}$
where $\mathbf{e}_{u}$ and $\mathbf{e}_{v}$ are standard basis vectors (vectors of zeros with a single 1 at position $u$ or $v$, respectively). Using the associativity of matrix multiplication, this becomes:
(17) $(\Delta B_{i}^{t})_{uv}\cdot\mathbf{e}_{u}\otimes(\mathbf{e}_{v}^{T}A_{i}^{t-1})$
The term $\mathbf{e}_{v}^{T}A_{i}^{t-1}$ is a vector product that selects the $v$-th row of the matrix $A_{i}^{t-1}$, which we denote as $(A_{i}^{t-1})_{v,:}$. Thus, the contribution of the single element $(\Delta B_{i}^{t})_{uv}$ to the final update matrix is itself a rank-1 matrix:
(18) $\text{Contribution Matrix}=(\Delta B_{i}^{t})_{uv}\cdot(\mathbf{e}_{u}\otimes(A_{i}^{t-1})_{v,:})$
This matrix is zero everywhere except for its $u$-th row, which is the $v$-th row of $A_{i}^{t-1}$ scaled by the scalar $(\Delta B_{i}^{t})_{uv}$.
A natural way to measure the magnitude, or ”importance”, of this contribution matrix is to compute its Frobenius norm ($\|\cdot\|_{F}$). The Frobenius norm of a scaled outer product of two vectors $\mathbf{x}$ and $\mathbf{y}$ is $|c|\cdot\|\mathbf{x}\|_{2}\cdot\|\mathbf{y}\|_{2}$. In our case, $c=(\Delta B_{i}^{t})_{uv}$, $\mathbf{x}=\mathbf{e}_{u}$, and $\mathbf{y}=(A_{i}^{t-1})_{v,:}$. Since $\|\mathbf{e}_{u}\|_{2}=1$, the norm is:
(19) $I((\Delta B_{i}^{t})_{uv})$ $=\left\|(\Delta B_{i}^{t})_{uv}\cdot(\mathbf{e}_{u}\otimes(A_{i}^{t-1})_{v,:})\right\|_{F}$
$=|(\Delta B_{i}^{t})_{uv}|\cdot\|\mathbf{e}_{u}\|_{2}\cdot\|(A_{i}^{t-1})_{v,:}\|_{2}$
$=|(\Delta B_{i}^{t})_{uv}|\cdot\|(A_{i}^{t-1})_{v,:}^{T}\|_{2}$
This is precisely the importance score for an element in $\Delta B_{i}^{t}$ as defined in Equation 6. A symmetric derivation for the term $B_{i}^{t}\Delta A_{i}^{t}$ shows that the importance of $(\Delta A_{i}^{t})_{uv}$ is its magnitude multiplied by the L2-norm of the $u$-th column of $B_{i}^{t}$. This formulation provides a principled way to measure a parameter’s true structural contribution to the final update, moving beyond its simple magnitude.
Appendix C Description of Baselines
Our baselines span from foundational FL algorithms to advanced techniques specifically designed for communication-efficient and effective federated fine-tuning of large language models.
FedAvg (McMahan et al., 2017) is the first algorithm in federated learning. It establishes a simple yet effective paradigm for collaborative model training. In each communication round, clients download current global model parameters. Each client then performs multiple steps of local training on its private dataset. The resulting updated model weights are sent back to the server, which aggregates them by computing a weighted average to produce the global model for the subsequent round. FedAvg serves as the fundamental baseline for performance and communication cost.
FedProx (Li et al., 2020) is a key extension to FedAvg designed to address the challenge of statistical heterogeneity (non-IID data) across clients. Non-IID data can cause local client models to drift significantly from the global consensus, leading to convergence instability. FedProx mitigates this by introducing a proximal term to the local objective function on each client. This term regularizes the local training process by penalizing the squared Euclidean distance between the local model weights and the global model weights, thereby constraining local updates and promoting more stable convergence.
FFA-LoRA (Sun et al., 2024) is a parameter-efficient fine-tuning method tailored for the federated setting, focusing on reducing communication overhead and accurate aggregation. It utilizes the Low-Rank Adaptation (LoRA) strategy. To achieve communication efficiency, FFA-LoRA freezes the randomly initialized LoRA $A$ matrices across all clients and throughout the training process. Consequently, clients only need to train and communicate the $B$ matrices, drastically reducing the number of parameters transmitted in each round.
FlexLoRA (Bai et al., 2024) proposes a more sophisticated aggregation mechanism for federated LoRA fine-tuning. It posits that directly averaging the low-rank matrices from different clients may be suboptimal. Instead, clients upload their locally trained LoRA matrices. The server first reconstructs the full-rank weight updates for each client ($B_{i}A_{i}$) and averages them. Subsequently, the server performs Singular Value Decomposition (SVD) to find a new, optimal low-rank approximation, which is then decomposed into new $B$ and $A$ matrices for client downloading.
FedICU (Liao et al., [n. d.]) is an advanced FL method that decomposes the global model into consensus and divergence components to handle heterogeneity better. The core mechanism is a consensus-divergence splitting performed on the server-side during aggregation. Consensus components and divergence components of all clients are aggregated separately. FEdICU further proposes the selective updating strategy that uploads only more specialized parameters for aggregation.
Fed-DARE addresses the communication bottleneck in FL by integrating the DARE (Drop and Rescale) technique (Yu et al., 2024). In this approach, clients sparsify their local LoRA updates by randomly dropping parameters. The remaining non-zero elements are then rescaled to preserve the update’s expected magnitude, resulting in highly sparse matrices that significantly reduce the communication payload. We adopt a sparsity ratio of 0.9 as recommended by DARE. As shown in Figure 2, this random-based approach outperforms magnitude-based sparsification, while higher sparsity ratios lead to performance degradation. This justifies our choice and the exclusion of magnitude-based methods as a baseline.
Table 8. Average In-domain Performance Comparison
Performance
FedAvg 33.9
   with Random sparsification (i.e., Fed-DARE) 33.57
   with Magnitude sparsification 31.61
   with Importance-aware sparsification (Our proposed) 34.04
Table 9. Communication cost comparison per round (MB). Costs include both parameters and bitmaps where applicable.
Method Upload (MB) Download (MB) Total (MB)
FedAvg / FedProx / FlexLoRA 371 371 742
FFA-LoRA 185.5 185.5 371
FedICU $\approx$197 371 $\approx$568
Fed-DARE $\approx$48.7 $\approx$141.5 $\approx$190
Ours (FedSRD / FedSRD-e) $\approx$31.0 $\approx$43.0 $\approx$74.0
Appendix D Communication Cost Analysis
We analyze the communication cost per round for each method, using a Llama3.2-3B model fine-tuned with LoRA (rank $r=64$) as a case study. The model contains 97,255,424 LoRA parameters stored in float32 format. The baseline payload for transmitting the full, dense parameter set in one direction (upload or download) is $\frac{97,255,424\times 4\text{ bytes}}{1024^{2}\text{ bytes/MB}}=371$ MB. For sparse methods, an additional bitmap is required to encode the positions of non-zero parameters, costing $\frac{97,255,424\text{ bits}}{8\text{ bits/byte}\times 1024^{2}\text{ bytes/MB}}\approx 11.6$ MB for the full set.
Table 9 details the communication costs for each method. Baseline approaches like FedAvg, FedProx, and FlexLoRA require a full round-trip transmission, totaling $742$ MB. FFA-LoRA halves this cost by communicating only the LoRA $B$ matrices. FedICU uses a selective upload ($\approx$50% sparsity) and a dense download, for a total of $568$ MB. Fed-DARE, with a 90% sparsity on uploads and an aggregated slightly more than 60% sparsity on downloads, lowers the cost to $190$ MB. Our proposed method achieves the highest communication efficiency. Client-side importance-aware pruning results in a lean 31 MB upload. For download, the server transmits only half of the parameters (alternating between LoRA $A$ and $B$ matrices) with 80% sparsity, leading to a download cost of just $43$ MB. This culminates in a total round-trip cost of only 74 MB, representing a 90% reduction compared to FedAvg. The overhead from bitmaps can be further mitigated using lossless compression techniques like Run-Length Encoding (Golomb, 1966).
Appendix E Importance-aware Sparsification Suits Well for FedAvg
To validate the superiority of our proposed importance-aware sparsification over others, we plug and compare them on FedAvg in Table 8. The results show that our importance-aware sparsification is more effective in the FL system.
Appendix F Justification for Equation 7
Equation 7 proposes a data-driven approach to dynamically determine the layer-wise sparsity ratio, $\rho$, for the LoRA update matrices $\Delta A_{i}^{t}$ and $\Delta B_{i}^{t}$. The formulation is designed to adapt the pruning aggressiveness based on the intrinsic characteristics of each layer’s importance score distribution, rather than relying on a uniform, manually-tuned sparsity level. The rationale behind each component of the equation is as follows:
Kurtosis as a Heuristic for Prunability: The core of our adaptive mechanism lies in using the kurtosis of the importance score distribution, $I(\cdot)$, as a quantitative indicator of a layer’s ”prunability.” Kurtosis measures the ”tailedness” of a distribution. A high kurtosis (a leptokurtic distribution) signifies that the importance scores are heavily concentrated on a small subset of parameters. This indicates a high degree of functional specialization and redundancy within the layer, making it an ideal candidate for aggressive pruning with minimal impact on performance. Conversely, a low kurtosis (a platykurtic distribution) implies that importance is more evenly diffused across many parameters. In this scenario, aggressive pruning would be detrimental, as it would likely remove parameters that make a non-trivial contribution. Therefore, a more conservative sparsity ratio is needed. By linking the sparsity ratio directly to kurtosis, our method automatically allocates a higher pruning budget to layers that can tolerate it and protects those that cannot.
Logarithmic Scaling: The raw values of kurtosis can span a wide range across different layers and training steps. We employ a logarithmic function to compress the wide range of kurtosis values into a more stable and manageable scale. This ensures that while a higher kurtosis leads to a higher sparsity ratio, the relationship is less sensitive to outlier kurtosis values, promoting more stable and robust pruning behavior throughout the process.
Modulation Coefficient: This coefficient scales the log-kurtosis term, modulating the final sparsity ratio. Empirically, we observe kurtosis values mostly ranging from 1 to 12 (with extremes up to 40). We selected a coefficient of 0.1. This choice is well-justified as it maps the typical kurtosis range to a moderate adjustment interval of approximately [0, 0.25]. This ensures the kurtosis contributes meaningfully to the sparsity calculation without introducing instability, making 0.1 a reasoned and effective choice.
Base Sparsity Ratio ($\alpha$): The hyperparameter $\alpha$ functions as a global base sparsity ratio. It establishes a baseline sparse level for the entire model, providing a direct and interpretable control over the overall sparsification ratio. The kurtosis-based term then acts as a dynamic, layer-specific offset to this baseline.
In summary, Equation 7 implies the insight that the distribution of parameter importance reveals its tolerance to sparsifying. It begins with a global sparsity baseline ($\alpha$), which is then intelligently adjusted for each layer using a logarithmically scaled measure of its importance concentration (kurtosis), thereby achieving a robust, adaptive, and effective sparsity allocation strategy. We leave other possible designs for adaptive determination of sparsity ratio for future work.
Appendix G Magnitude of $B^{t-1},A^{t-1}$ and $\Delta B^{t},\Delta A^{t}$
In the Taylor expansion in Equation 11, the product term $\Delta B^{t}\Delta A^{t}$ is treated as a negligible higher-order term. To validate this assumption, we empirically analyze the magnitudes during the federated fine-tuning process. The average relative magnitude for matrix $A$, denoted as $\frac{\|\Delta A^{t}\|}{\|A^{t-1}\|}$, is less than 3%, while for matrix $B$, $\frac{\|\Delta B^{t}\|}{\|B^{t-1}\|}$, it is less than 10%. Their product constitutes a higher-order term that can be safely omitted. This conclusion is further reinforced by the observed convergence trend, where both relative ratios systematically decrease as the number of federated rounds increases.
Appendix H Performance of Base Models and Locally Trained Models
We test the performance of base models (Llama3.2-3B and Qwen2-7B) on our benchmarks. We also conduct local training on each client (e.g., Code indicates the model is locally trained on the client from the code domain) and test the locally trained models on each client on our benchmarks. The results are reported in Table 10.
Table 10. Performance of base models and local training
Llama3.2-3B Qwen2-7B
Model In-Domain OOD In-Domain OOD
Base 16.33 14.90 44.79 38.83
Code 29.74 24.19 48.93 41.84
Medical 25.49 23.65 52.11 40.92
Finance 27.73 23.33 49.44 35.29
Math 19.25 18.80 43.83 34.16

### v2

@@
 The structural redundancy within LoRA parameters not only incurs a heavy communication burden but also introduces conflicts when aggregating client updates.
-To address this, we propose FedSRD, a Sparsify-Reconstruct-Decompose framework designed for communication-efficient FL.
+To address this, we propose FedSRD, a Sparsify-Reconstruct-Decompose framework designed for communication-efficient federated LLMs fine-tuning.
 We first introduce an importance-aware sparsification method that preserves the structural integrity of LoRA updates to reduce the uploaded parameter count.
@@
 However, realizing their full potential for specialized downstream applications requires fine-tuning on diverse, high-quality, domain-specific data.
-Such data is often proprietary and distributed across numerous private sources, resulting in fragmented ”data silos.” As centralizing this sensitive data is frequently impractical due to privacy regulations and cost constraints, Federated Learning (FL) offers a compelling solution, enabling multiple clients to collaboratively fine-tune a global shared LLM without direct data exchange.
-This paradigm is central to building the next generation of intelligent, decentralized Web applications.
+Such data is often proprietary and distributed across numerous private sources, resulting in fragmented ”data silos”.
+As centralizing this sensitive data is frequently impractical due to privacy regulations and cost constraints, Federated Learning (FL) offers a compelling solution, enabling multiple clients to collaboratively fine-tune a global shared LLM without direct data exchange.
+This paradigm is central to building the next generation of intelligent, decentralized Web applications (Sani et al., [n.
+d.]; Wu et al., 2025).
 While FL provides a privacy-preserving solution, fine-tuning billion-parameter LLMs across distributed clients remains computationally and communicatively prohibitive for individual participants.
@@
 While sparsification is a common strategy for reducing communication, conventional methods are ill-adapted for federated LoRA fine-tuning on non-IID data.
-As shown in Figure 2, naïve approaches like magnitude-based pruning (Aji and Heafield, 2017; Yadav et al., 2023; Deep et al., 2024) disrupt the critical structural relationship between the LoRA matrices $A$ and $B$, while random sparsification (Yu et al., 2024) also leads to performance degradation at high sparsity ratios.
+As shown in Figure 2, naïve approaches like magnitude-based pruning (Aji and Heafield, 2017) disrupt the critical structural relationship between the LoRA matrices $A$ and $B$, while random sparsification (Yu et al., 2024) also leads to performance degradation at high sparsity ratios.
 Crucially, both methods fail to preserve client-specific knowledge and exacerbate conflicts during server-side aggregation under data heterogeneity.
@@
 Both random and magnitude-based sparsification suffer performance decline at high sparsity ratios, motivating the need for a more advanced, structure-aware strategy.
-To address this question, we introduce FedSRD, a novel Sparsify-Reconstruct-Decompose framework for communication-efficient federated LLM fine-tuning.
+To address this question, we introduce FedSRD, a novel Sparsify-Reconstruct-Decompose framework for communication-efficient federated LLMs fine-tuning.
 Our framework is designed to maintain high model performance through three sequential operations.
@@
 For instance, FFA-LoRA (Sun et al., 2024) halves the communication cost by transmitting only the $B$ matrices.
-Building on this, FedSVD (Lee et al., 2025) iteratively re-initializes LoRA $A$ matrices.
 Fed-SB (Singhal et al., 2025) freezes the LoRA matrices entirely and instead inserts a small, trainable square matrix for aggregation.
@@
 Motivation
-After aggregating in the full-rank space, the server must decompose the result back into low-rank matrices for the clients.
-Previous works (Bai et al., 2024; Lee et al., 2025) often use Singular Value Decomposition (SVD) to construct new dense matrices $B^{t}$ and $A^{t}$.
-However, transmitting these dense matrices would create a severe download bottleneck, negating the benefit of a sparse upload.
-Naively sparsifying the full state matrices $B^{t}$ and $A^{t}$ would damage the learned model, and one cannot simply compute an update as $B^{t}-B^{t-1}$ because it does not correctly approximate the global update ($B^{t}A^{t}-B^{t-1}A^{t-1}\neq(B^{t}-B^{t-1})A^{t-1}$).
-This motivates a novel decomposition method that can produce a sparse update for download.
+Previous works (Bai et al., 2024; Lee et al., 2025) often use Singular Value Decomposition (SVD) to decompose aggregated weights into new dense matrices $B^{t}$ and $A^{t}$.
+But transmitting them creates a download bottleneck.
+Naively sparsifying the full state matrices $B^{t}$ and $A^{t}$ would damage the learned model.
+Computing the update as a direct difference, $B^{t}-B^{t-1}$, is also mathematically flawed.
+This is because SVD is rotation-invariant and decomposition is not unique.
+Even a small global update can induce an arbitrary rotation of the low-rank basis between rounds, creating large, dense updates that do not reflect the true minimal updates.
+Thus, they are highly unsuitable for sparsification.
+These motivate the need for a novel decomposition capable of producing a stable update for sparse downloads.
 Our goal is to find sparse updates $\Delta B^{t}$ and $\Delta A^{t}$ that approximate the change from the previous global state $W^{t-1}$ to the new one $W^{t}$.
@@
 For our methods, while the full FedSRD requires more time for its $r$-rank SVD projection, our efficient variant, FedSRD-e, introduces negligible additional computation compared to FedAvg, while exhibiting performance second only to the full FedSRD.
-This result highlights that our FedSRD-e variant achieves a Pareto-optimal balance across model performance, communication efficiency, and time consumption, making it a highly practical and scalable solution for real-world federated LLM fine-tuning.
+This result highlights that our FedSRD-e variant achieves a Pareto-optimal balance across model performance, communication efficiency, and time consumption, making it a highly practical and scalable solution for real-world federated LLMs fine-tuning.
 Table 7. Computation time per round for different methods.
@@
 In this paper, we address the critical challenge of communication overhead in federated LoRA fine-tuning for LLMs.
-We propose FedSRD, a novel Sparsify-Reconstruct-Decompose framework for communication-efficient FL.
+We propose FedSRD, a novel Sparsify-Reconstruct-Decompose framework for communication-efficient federated LLMs fine-tuning.
 Its core components include a structure-aware, importance-based sparsification strategy on the client side; a conflict-mitigating reconstruction and aggregation step in the full-rank space on the server side; and a novel alternating decomposition method to produce sparse global updates for efficient download.

### v3

@@
 Abstract.
-The current paradigm of training large language models (LLMs) on publicly available Web data is becoming unsustainable, with high-quality data sources in specialized domains nearing exhaustion.
-Federated Learning (FL) emerges as a practical solution for the next generation of AI on a decentralized Web, enabling privacy-preserving collaborative fine-tuning by leveraging private data distributed across a global client base.
-While Low-Rank Adaptation (LoRA) is the standard for efficient fine-tuning, its application in federated settings presents a critical challenge: communication overhead remains a significant bottleneck across the Web’s heterogeneous network conditions.
-The structural redundancy within LoRA parameters not only incurs a heavy communication burden but also introduces conflicts when aggregating client updates.
-To address this, we propose FedSRD, a Sparsify-Reconstruct-Decompose framework designed for communication-efficient federated LLMs fine-tuning.
-We first introduce an importance-aware sparsification method that preserves the structural integrity of LoRA updates to reduce the uploaded parameter count.
-The server then reconstructs and aggregates these updates in a full-rank space to mitigate conflicts.
-Finally, it decomposes the global update into a sparse low-rank format for broadcast, ensuring a symmetrically efficient cycle.
+The current paradigm of training large language models (LLMs) on public available Web data is becoming unsustainable as high-quality data sources in specialized domains near exhaustion.
+Federated Learning (FL) emerges as a practical solution for the next generation of AI on a decentralized Web, enabling privacy-preserving collaborative fine-tuning on decentralized private data.
+While Low-Rank Adaptation (LoRA) is standard for efficient fine-tuning, its federated application faces a critical bottleneck: communication overhead under heterogeneous network conditions.
+Structural redundancy in LoRA parameters increases communication costs and causes aggregation conflicts.
+To address this, we propose FedSRD, a Sparsify-Reconstruct-Decompose framework for communication-efficient federated LLM fine-tuning.
+We introduce importance-aware sparsification to reduce the upload parameter count while preserving the structural integrity of LoRA updates.
+The server aggregates updates in full-rank space to mitigate conflicts, then decomposes the global update into a sparse low-rank format for broadcast, ensuring a symmetrically efficient cycle.
 We also propose an efficient variant, FedSRD-e, to reduce computational overhead.
-Experimental results on 10 benchmarks demonstrate that our framework significantly reduces communication costs by up to 90% while even improving model performance on heterogeneous client data.
-Federated Learning, Large Language Models, Communication, LoRA, Fine-tuning
-copyright: nonedoi: XXXXXXX.XXXXXXXccs: Computing methodologies Cooperation and coordination
+Experiments on 10 benchmarks show our framework significantly reduces communication costs by up to 90% while improving performance on heterogeneous client data.
+Federated Learning; Large Language Models; Communication; LoRA
+journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emirates.booktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesisbn: 979-8-4007-2307-0/2026/04doi: 10.1145/3774904.3792144ccs: Computing methodologies Cooperation and coordination
 1. Introduction
@@
 As centralizing this sensitive data is frequently impractical due to privacy regulations and cost constraints, Federated Learning (FL) offers a compelling solution, enabling multiple clients to collaboratively fine-tune a global shared LLM without direct data exchange.
-This paradigm is central to building the next generation of intelligent, decentralized Web applications (Sani et al., [n.
-d.]; Wu et al., 2025).
-While FL provides a privacy-preserving solution, fine-tuning billion-parameter LLMs across distributed clients remains computationally and communicatively prohibitive for individual participants.
+This paradigm is central to building the next generation of intelligent, decentralized Web applications (Sani et al., ; Wu et al., 2025).
+While FL provides a privacy-preserving solution, fine-tuning billion-parameter LLMs across distributed clients remains computationally and communicatively prohibitive.
 To overcome this, Parameter-Efficient Fine-Tuning (PEFT) methods, particularly Low-Rank Adaptation (LoRA) (Hu et al., 2022), have become the standard (Ding et al., 2023; Kuang et al., 2024; Ye et al., 2024).
-As illustrated in Figure 1, LoRA significantly reduces the training burden by freezing the pre-trained model weights and optimizing only a small set of trainable, low-rank matrices.
+As illustrated in Figure 1, LoRA significantly reduces the training burden by optimizing only a small set of trainable, low-rank matrices.
 In a federated setting, clients train and communicate only these LoRA parameters, achieving performance comparable to full-parameter fine-tuning at a fraction of the resource cost.
-Despite the efficiency of LoRA, the iterative communication of its parameters remains a substantial overhead.
-This becomes a critical bottleneck when deployed across the highly heterogeneous and unpredictable network conditions characteristic of the global Web and mobile systems (Dorfman et al., 2023; Mundt, 2004).
+Despite the efficiency, the iterative communication of LoRA parameters remains a substantial overhead.
+This becomes a critical bottleneck when deployed across the heterogeneous and unpredictable network conditions characteristic of the global Web and mobile systems (Dorfman et al., 2023; Mundt, 2004).
 High communication costs can exclude clients with limited bandwidth, such as those in rural areas or on mobile devices, undermining data diversity and posing fairness and inclusivity challenges for the global Web ecosystem.
 While sparsification is a common strategy for reducing communication, conventional methods are ill-adapted for federated LoRA fine-tuning on non-IID data.
-As shown in Figure 2, naïve approaches like magnitude-based pruning (Aji and Heafield, 2017) disrupt the critical structural relationship between the LoRA matrices $A$ and $B$, while random sparsification (Yu et al., 2024) also leads to performance degradation at high sparsity ratios.
+As shown in Figure 2, naïve approaches like magnitude-based pruning (Aji and Heafield, 2017) disrupt the structural relationship between LoRA matrices $A$ and $B$, while random sparsification (Yu et al., 2024) also degrades at high sparsity ratios.
 Crucially, both methods fail to preserve client-specific knowledge and exacerbate conflicts during server-side aggregation under data heterogeneity.
 This results in an undesirable trade-off: reduced communication at the expense of significant performance loss.
-This leads us to our central research question:
+This leads to our central research question:
 How can we significantly reduce communication costs in federated LLMs fine-tuning without sacrificing model performance under heterogeneous client data?
@@
 To address this question, we introduce FedSRD, a novel Sparsify-Reconstruct-Decompose framework for communication-efficient federated LLMs fine-tuning.
-Our framework is designed to maintain high model performance through three sequential operations.
+Our framework ensures high model performance through three sequential operations.
 First, on the client side, we develop an importance-aware Sparsification strategy.
-Instead of relying on parameter magnitude, we calculate the importance scores that quantify the contribution of each element in the update matrices ($\Delta B$ and $\Delta A$) to the overall low-rank update ($\Delta W$).
-This effectively preserves the critical structural interplay between the LoRA matrices.
-Furthermore, recognizing that different layers in an LLM serve distinct functions, FedSRD adaptively determines the sparsity ratio for each layer based on the kurtosis of its importance score distribution.
-Aggressive pruning is applied where high importance is concentrated in a few parameters, while conservative pruning is adopted where importance is more diffuse.
+Instead of relying on parameter magnitude, we calculate the importance scores that quantify the contribution of each element in update matrices ($\Delta B$ and $\Delta A$) to the overall update ($\Delta W$).
+This effectively preserves the critical structural interplay between LoRA matrices.
+Furthermore, recognizing that different layers serve distinct functions, we adaptively determine layer-wise sparsity ratios based on the kurtosis of its importance scores.
+Aggressive pruning is applied where high importance is concentrated.
 Second, the server executes the Reconstruction and Decomposition.
-Upon receiving sparse updates, the server first reconstructs each client’s full-rank weight update ($\Delta W$) before aggregation.
-This enables aggregation to occur in the full-rank parameter space, a process that is mathematically sound and effectively mitigates the update conflicts arising from non-IID data (Wang et al., 2024; Sun et al., 2024).
+Upon receiving sparse updates, the server first reconstructs each client’s full-rank update ($\Delta W$) before aggregation.
+This enables aggregation to occur in the full-rank space, which is mathematically sound and effectively mitigates the update conflicts from non-IID data (Wang et al., 2024; Sun et al., 2024).
 Finally, to create a symmetrically efficient communication cycle, the server performs a novel Decomposition of the aggregated global update back into a sparse, low-rank format for clients to download.
 Using a Taylor expansion-based approximation, we employ an alternating scheme that solves for and transmits only one sparse low-rank matrix per round.
-Furthermore, for scenarios where server-side computation is the bottleneck (e.g., on edge servers), we propose an efficient variant, FedSRD-e, which bypasses the computationally intensive SVD projection during decomposition.
-We demonstrate the effectiveness of our framework through extensive experiments that simulate practical non-IID scenarios by finetuning an LLM with clients from diverse specialist domains.
+Furthermore, We propose FedSRD-e, a computational-efficient variant that bypasses intensive SVD operations for resource-constrained servers.
+We demonstrate the effectiveness of FedSRD through extensive experiments that simulate practical non-IID scenarios by finetuning an LLM with clients from diverse specialist domains.
 Our evaluation on 8 in-domain and 2 out-of-domain benchmarks shows that FedSRD reduces both upload and download communication by 90% while achieving even better performance.
@@
 2.1. Parameter-Efficient Fine-Tuning
-The large scale of modern LLMs makes the full-parameter fine-tuning computationally prohibitive in many resource-constrained environments, such as on mobile or edge devices.
-To address this, Parameter-Efficient Fine-Tuning (PEFT) methods (Ding et al., 2023) have been developed, which freeze most of a pre-trained model’s parameters and only update a small part.
-Current approaches include adapter-based methods (He et al., 2021; Edalati et al., 2025; Houlsby et al., 2019; Hu et al., 2022) that insert small, trainable modules between model layers, and prompting methods (Petrov et al., 2023; Lester et al., 2021), such as prefix-tuning (Li and Liang, 2021), that add learnable tokens to the input sequence.
+The large scale of modern LLMs makes the full-parameter fine-tuning computationally prohibitive in resource-constrained environments.
+To address this, Parameter-Efficient Fine-Tuning (PEFT) methods (Ding et al., 2023) freeze most of model parameters and only update a small part.
+Current approaches include adapter-based methods (He et al., 2021; Edalati et al., 2025; Houlsby et al., 2019; Hu et al., 2022) that insert small, trainable modules in models, and prompting methods (Petrov et al., 2023; Lester et al., 2021), such as prefix-tuning (Li and Liang, 2021), that add learnable tokens to the input sequence.
 Among these, Low-Rank Adaptation (LoRA) (Hu et al., 2022) has emerged as a dominant and effective approach.
-The key insight of LoRA is that weight updates during fine-tuning possess a low intrinsic rank.
-LoRA represents the update for each target weight matrix as the product of two smaller, trainable low-rank matrices, significantly reducing the number of trainable parameters and associated memory footprint while achieving comparable performance.
+The key insight of LoRA is that weight updates in fine-tuning possess a low intrinsic rank.
+LoRA represents each weight matrix update as the product of two smaller, trainable low-rank matrices, significantly reducing the number of trainable parameters and memory footprint while achieving comparable performance.
 2.2. Federated LLMs Fine-tuning
@@
 FlexLoRA (Bai et al., 2024) proposes to aggregate the true model updates ($BA$) on the server and then employs Singular Value Decomposition (SVD) to derive new global low-rank matrices for clients.
-FedICU (Liao et al., [n.
-d.]) disentangles client updates into consensus and divergence components, which are then aggregated separately.
+FedICU (Liao et al., ) disentangles client updates into consensus and divergence components, which are then aggregated separately.
 2.3. Communication-Efficient Federated LLMs Finetuning
@@
 4. Proposed Method
-In this section, we introduce FedSRD, a Sparsify-Reconstruct-Decompose framework for federated LoRA fine-tuning, illustrated in Figure 3.
-FedSRD establishes a symmetrically efficient communication cycle by transmitting only sparse updates of the LoRA parameters.
+In this section, we introduce FedSRD, a noval Sparsify-Reconstruct-Decompose framework for federated LoRA fine-tuning, illustrated in Figure 3.
 At each communication round $t$, a client $i$ first computes its local LoRA updates $\Delta B_{i}^{t}$ and $\Delta A_{i}^{t}$, then uses our proposed importance-aware strategy to sparsify them for upload (Section 4.1).
@@
 Finally, to ensure an efficient download, the server decomposes the resulting global update into a single sparse low-rank matrix for broadcast using an alternating, Taylor approximation-based method (Section 4.3).
-The pseudo-code is provided in Appendix A.
 Figure 3.
@@
 A higher importance score indicates a greater contribution to the final weight update.
-We explain the rationale behind in Appendix B.
-We then prune elements with the lowest scores, transmitting only the most important portion of the parameters.
-Furthermore, recognizing that LoRA matrices in different layers play distinct roles (Luo et al., 2025), we propose to adaptively assign a sparsity ratio $\rho$ to each matrix.
-The core intuition is to use the kurtosis of the importance score distribution reveals a layer’s functional ”prunability”.
+We then prune elements with the lowest scores.
+We explain the rationale behind and prove that pruning based on this score minimizes the upper bound of the weight update approximation error $\|\Delta W_{i}^{t}\|_{F}$ in Appendix B.
+Furthermore, recognizing that LoRA matrices in different layers play distinct roles (Luo et al., 2025), we adaptively assign a sparsity ratio $\rho$ to each matrix.
+The core intuition is that the kurtosis of the importance score reveals a layer’s functional ”prunability”.
 A high kurtosis (a leptokurtic, heavy-tailed distribution) signifies that importance is concentrated in a few parameters, justifying an aggressive sparsity ratio.
@@
 The sparsity ratios are truncated if they exceed a pre-set upper bound to ensure validity.
-We provide a detailed explanation and justification of this design in Appendix F.
+We provide a detailed explanation and justification of this design in Appendix E.
 We use the adaptive sparsity ratios to prune $\Delta B_{i}^{t}$ and $\Delta A_{i}^{t}$:
@@
 Most existing federated learning methods like FedAvg (McMahan et al., 2017), when applied to LoRA, average the clients’ LoRA matrices independently.
-This aggregation is not mathematically sound and introduces unexpected noises because it multiplicative relationship between the matrices and fails to mitigate update conflicts arising from non-IID client data (Sun et al., 2024; Bai et al., 2024; Wang et al., 2024).
+This aggregation is not mathematically sound and introduces unexpected noises because its multiplicative relationship between the matrices and fails to mitigate update conflicts arising from non-IID client data (Sun et al., 2024; Bai et al., 2024; Wang et al., 2024).
 This motivates us to develop an advanced aggregation method.
@@
 $\approx B^{t-1}\Delta A^{t}+\Delta B^{t}A^{t-1},$
-where the higher-order term $\Delta B^{t}\Delta A^{t}$ is omitted (discussed in Appendix G).
+where the higher-order term $\Delta B^{t}\Delta A^{t}$ is omitted (discussed in Appendix F).
 To solve this underdetermined system, we alternate between setting one of the deltas to zero.
@@
 5.1.2. Baselines
-We compare our framework FedSRD and FedSRD-e with 6 state-of-the-art federated learning methods: (1) FedAvg (McMahan et al., 2017), (2) FedProx (Li et al., 2020), (3) FFA-LoRA (Sun et al., 2024), which freezes LoRA $A$ matrcies to ensure efficient communication and effective aggregation, (4) FlexLoRA (Bai et al., 2024), which aggregates in full-rank space and uses SVD for download, (5) FedICU (Liao et al., [n.
-d.]), which selectively updates on global model and perform consensus-divergence splitting in the server-side aggregatio, (6) Fed-DARE, which applies random sparsification with rescaling (DARE (Yu et al., 2024)) to LoRA updates for upload.
-A detailed description of baselines is in Appendix C.
+We compare our framework FedSRD and FedSRD-e with 6 state-of-the-art federated learning methods: (1) FedAvg (McMahan et al., 2017), (2) FedProx (Li et al., 2020), (3) FFA-LoRA (Sun et al., 2024), which freezes LoRA $A$ matrcies to ensure efficient communication and effective aggregation, (4) FlexLoRA (Bai et al., 2024), which aggregates in full-rank space and uses SVD for download, (5) FedICU (Liao et al., ), which selectively updates on global model and perform consensus-divergence splitting in the server-side aggregatio, (6) Fed-DARE, which applies random sparsification with rescaling (DARE (Yu et al., 2024)) to LoRA updates for upload.
 5.1.3. Benchmarks
@@
 Table 3.
-Comparison of model performance and per-round communication cost using Llama3.2-3B as the base model.
+Comparison of model performance and per-round communication cost.
+The top section uses Llama3.2-3B and the bottom section uses Qwen2-7B.
 OOD represents Out-of-Domain.
-The best and second-best results are highlighted in bold and underline, respectively.
-Method
-Comm.
-Cost
-Average In-Domain Out-of-Domain
+Best and second-best results are bolded and underlined.
+Method Comm. Cost (MB) Average In-Domain Out-of-Domain
 In-Domain OOD Code Medical Finance Math AGIEval Law
 HumanEval MBPP MedQA MedMCQA FinEval FinanceIQ GSM8K MATH
+Base Model: Llama3.2-3B
 FedAvg 742 33.90 24.36 26.85 47.86 40.45 41.53 41.96 38.08 28.96 5.56 22.47 26.24
@@
 FedSRD 74 35.12 26.17 31.10 49.42 41.95 41.62 43.44 39.13 28.96 5.36 25.54 26.79
-Table 4.
-Comparison of model performance and per-round communication cost using Qwen2-7B as the base model.
-OOD represents Out-of-Domain.
-The best and second-best results are highlighted in bold and underline, respectively.
-Method
-Comm.
-Cost
-Average In-Domain Out-of-Domain
-In-Domain OOD Code Medical Finance Math AGIEval Law
-HumanEval MBPP MedQA MedMCQA FinEval FinanceIQ GSM8K MATH
+Base Model: Qwen2-7B
 FedAvg 616 57.04 41.02 59.15 37.35 61.67 75.23 65.16 49.04 64.29 44.46 48.25 33.79
@@
 5.2.1. Performance Comparison
-The main results, reported in Table 3 (Llama3.2-3B) and Table 4 (Qwen2-7B), comprehensively demonstrate the effectiveness and efficiency of our FedSRD framework.
-We analyze performance from three key perspectives.
+The main results are reported in Table 3. We analyze the results from three key perspectives.
 In-Domain Performance.
@@
 As shown in Table 3, with the Llama3.2-3B model, FedSRD and FedSRD-e achieve the highest average in-domain performance score, outperforming all baseline methods.
-This trend is further confirmed with the larger Qwen2-7B model in Table 4, where FedSRD and FedSRD-e again lead with top performance, respectively.
+This trend is further confirmed with the larger Qwen2-7B model, where FedSRD and FedSRD-e again lead with top performance, respectively.
 While some baselines may achieve peak performance on specific benchmarks, they often falter in other domains.
@@
 A standout advantage of our framework is its remarkable communication efficiency.
-FedSRD and FedSRD-e reduce the per-round communication overhead to just 74 MB for Llama3.2-3B and 61 MB for Qwen2-7B.
-This represents a reduction of $\approx$ 90% compared to standard methods like FedAvg (742 MB) and advanced methods like FlexLoRA (742 MB).
-This reduction is achieved by our proposed importance-aware sparsification for uploads and alternative decomposition for downloads (see Appendix D for a detailed analysis).
-Crucially, this significant gain in efficiency does not come at the cost of performance.
+FedSRD and FedSRD-e reduce the per-round communication overhead to just 74 MB for Llama3.2-3B and 61 MB for Qwen2-7B, an approximate 90% reduction compared to standard methods like FedAvg (742 MB) and advanced methods like FlexLoRA (742 MB).
+This is achieved by our proposed importance-aware sparsification for uploads and alternative decomposition for downloads (see Appendix C for a detailed analysis).
+Crucially, this significant gain in efficiency does not compromise performance.
 On the contrary, FedSRD and FedSRD-e achieve higher performance than baselines, directly answering our central research question.
@@
 Out-of-Domain Performance.
-Our methods also demonstrate strong performance on out-of-domain (OOD) benchmarks, indicating effective mitigation of catastrophic forgetting.
-For Llama3.2-3B, FedSRD and FedSRD-e achieve the leading OOD performance.
-For Qwen2-7B, they also maintain competitive performance.
-Although FFA-LoRA also exhibits strong OOD retention, we attribute this to its mechanism of freezing the LoRA $A$ matrices.
+Our methods also show strong performance on out-of-domain (OOD) benchmarks, effectively mitigating catastrophic forgetting.
+FedSRD and FedSRD-e achieve leading results on Llama3.2-3B and remain competitive on Qwen2-7B.
+Although FFA-LoRA exhibits strong OOD retention, we attribute this to its mechanism of freezing the LoRA $A$ matrices.
 which inherently constrains the capacity and hinders new knowledge acquisition in favor of preserving prior information.
-In contrast, our methods provide a more balanced and consistently high level of performance across both in-domain and out-of-domain evaluations.
+In contrast, our methods provide a more balanced and consistently high performance across both in-domain and out-of-domain evaluations.
 Further, we posit that the SVD projection in FedSRD, by seeking a low-rank consensus for in-domain tasks, may inadvertently filter out some general knowledge from the base model, which could account for the superior OOD performance of FedSRD-e.
@@
 These results confirm that each component plays an indispensable and synergistic role in achieving the dual objectives of high model performance and extreme communication efficiency.
-Additionally, we show in Appendix E that our importance-aware sparsification also works as a more effective plugin to enhance conventional methods like FedAvg.
+Additionally, we show in Appendix D that our importance-aware sparsification also works as a more effective plugin to enhance conventional methods like FedAvg.
 Table 4.
@@
 5.3. Computation Time Analysis
-We measured the wall-clock time required for a single FL round in our simulation environment for both the Llama3.2-3B and Qwen2-7B models, with results reported in Table 7.
+We report the wall-clock time required for a single FL round in our simulation environment for both the Llama3.2-3B and Qwen2-7B in Table 6.
 Among the baselines, FedAvg provides a baseline for time consumption.
@@
 Our work presents a practical and high-performing solution that significantly lowers the barrier for scalable, collaborative AI development on the decentralized Web.
+Acknowledgements.
+This work was supported by the National Key R&D Program of China under Grant No.
+2022YFB2703301.
+This work was also supported by the Science and Technology Development Program of Two Districts in Xinjiang, China under Grant No.
+2024LQ03004.
 Appendix A Pseudo-Code of FedSRD and FedSRD-e
@@
 1:Server Initializes: Global LoRA matrices $B^{0},A^{0}$. Let $\mathcal{C}$ be the set of all clients.
-2:Initial Synchronization: Each client $i\in\mathcal{C}$ sets its local state $B_{i}^{0}\leftarrow B^{0},A_{i}^{0}\leftarrow A^{0}$.
-3:for each communication round $t=1,2,\dots,T$ do
-4:  if $t>1$ then
-5:   Server broadcasts $\Delta B_{s}^{t-1}$ or $\Delta A_{s}^{t-1}$.
-6:   for each client $i\in\mathcal{C}$ in parallel do
-7:     Client updates local model: $B_{i}^{t-1}\leftarrow B_{i}^{t-2}+\Delta B_{s}^{t-1}$, $A_{i}^{t-1}\leftarrow A_{i}^{t-2}+\Delta A_{s}^{t-1}$.
-8:  for each client $i\in\mathcal{C}$ in parallel do
-9:   $\Delta B_{i,s}^{t},\Delta A_{i,s}^{t}\leftarrow\text{ClientTraining}(i)$
-10:  Server-Side Reconstruction and Aggregation:
-11:  for each client $i\in\mathcal{C}$ do
-12:   Client reconstructs its new state $B_{i}^{t},A_{i}^{t}$ using Eq. 4
-13:   Server computes client’s full-rank matrix: $W_{i}^{t}\leftarrow B_{i}^{t}A_{i}^{t}$.
-14:  Aggregate in full-rank space: $W^{t}\leftarrow\frac{1}{|\mathcal{C}|}\sum_{i}W_{i}^{t}$.
-15:  Server-Side Alternating Decomposition:
-16:  $W^{t}_{r}\leftarrow\text{SVD}_{r}(W^{t})$. $\triangleright$ FedSRD-e bypasses this SVD step.
-17:  Let $\Delta W^{t}_{update}=W^{t}_{r}-B^{t-1}A^{t-1}$. $\triangleright$ For FedSRD-e, use $W^{t}$ instead of $W^{t}_{r}$.
-18:  if $t$ is even then
-19:   $\Delta A^{t}\leftarrow(B^{t-1})^{\dagger}(\Delta W^{t}_{update})$
-20:   $\Delta B^{t}\leftarrow\mathbf{0}$
-21:  else
-22:   $\Delta B^{t}\leftarrow(\Delta W^{t}_{update})(A^{t-1})^{\dagger}$
-23:   $\Delta A^{t}\leftarrow\mathbf{0}$
-24:  Update global state on server: $B^{t}\leftarrow B^{t-1}+\Delta B^{t}$, $A^{t}\leftarrow A^{t-1}+\Delta A^{t}$.
-25:  Sparsify for download: $\Delta B_{s}^{t},\Delta A_{s}^{t}\leftarrow\text{DARE}(\Delta B^{t},\Delta A^{t})$.
-26:
-27:function ClientTraining(ii)
-28:  Fine-tune model on local data $\mathcal{D}_{i}$ to get new $B^{\prime}_{i},A^{\prime}_{i}$.
-29:  Compute local updates: $\Delta B_{i}^{t}\leftarrow B^{\prime}_{i}-B_{i}^{t-1}$, $\Delta A_{i}^{t}\leftarrow A^{\prime}_{i}-A_{i}^{t-1}$.
-30:  Importance-Aware Sparsification:
-31:  Calculate importance scores $I(\Delta B_{i}^{t}),I(\Delta A_{i}^{t})$ using Eq. 6.
-32:  Calculate adaptive sparsity ratios $\rho_{\Delta B_{i}^{t}},\rho_{\Delta A_{i}^{t}}$.
-33:  Prune updates to get sparse $\Delta B_{i,s}^{t},\Delta A_{i,s}^{t}$ adaptively.
-34:  return $\Delta B_{i,s}^{t},\Delta A_{i,s}^{t}$ to server.
+2:for each communication round $t=1,2,\dots,T$ do
+3:  if $t>1$ then
+4:   Server broadcasts $\Delta B_{s}^{t-1}$ or $\Delta A_{s}^{t-1}$.
+5:   for each client $i\in\mathcal{C}$ in parallel do
+6:    Client updates local model: $B_{i}^{t-1}\leftarrow B_{i}^{t-2}+\Delta B_{s}^{t-1}$, $A_{i}^{t-1}\leftarrow A_{i}^{t-2}+\Delta A_{s}^{t-1}$.
+7:   end for
+8:  end if
+9:  for each client $i\in\mathcal{C}$ in parallel do
+10:   $\Delta B_{i,s}^{t},\Delta A_{i,s}^{t}\leftarrow\text{ClientTraining}(i)$
+11:  end for
+12:  Server-Side Reconstruction and Aggregation:
+13:  for each client $i\in\mathcal{C}$ do
+14:   Client reconstructs its new state $B_{i}^{t},A_{i}^{t}$ using Eq. 4
+15:   Server computes client’s full-rank matrix: $W_{i}^{t}\leftarrow B_{i}^{t}A_{i}^{t}$.
+16:  end for
+17:  Aggregate in full-rank space: $W^{t}\leftarrow\frac{1}{|\mathcal{C}|}\sum_{i}W_{i}^{t}$.
+18:  Server-Side Alternating Decomposition:
+19:  $W^{t}_{r}\leftarrow\text{SVD}_{r}(W^{t})$. {FedSRD-e bypasses this SVD step.}
+20:  Let $\Delta W^{t}_{update}=W^{t}_{r}-B^{t-1}A^{t-1}$. {For FedSRD-e, use $W^{t}$ instead of $W^{t}_{r}$.}
+21:  if $t$ is even then
+22:   $\Delta A^{t}\leftarrow(B^{t-1})^{\dagger}(\Delta W^{t}_{update})$
+23:   $\Delta B^{t}\leftarrow\mathbf{0}$
+24:  else
+25:   $\Delta B^{t}\leftarrow(\Delta W^{t}_{update})(A^{t-1})^{\dagger}$
+26:   $\Delta A^{t}\leftarrow\mathbf{0}$
+27:  end if
+28:  Update global state: $B^{t}\leftarrow B^{t-1}+\Delta B^{t}$, $A^{t}\leftarrow A^{t-1}+\Delta A^{t}$.
+29:  Sparsify for download: $\Delta B_{s}^{t},\Delta A_{s}^{t}\leftarrow\text{DARE}(\Delta B^{t},\Delta A^{t})$.
+30:end for
+31:
+32:Function ClientTraining(ii):
+33:  Fine-tune model on local data $\mathcal{D}_{i}$ to get new $B^{\prime}_{i},A^{\prime}_{i}$.
+[…]
+35:  Importance-Aware Sparsification:
+36:  Calculate importance scores $I(\Delta B_{i}^{t}),I(\Delta A_{i}^{t})$ using Eq. 6.
+37:  Calculate adaptive sparsity ratios $\rho_{\Delta B_{i}^{t}},\rho_{\Delta A_{i}^{t}}$.
+38:  Prune updates to get sparse $\Delta B_{i,s}^{t},\Delta A_{i,s}^{t}$ adaptively.
+39:return $\Delta B_{i,s}^{t},\Delta A_{i,s}^{t}$ to server.
 Appendix B Rationale of the Importance Score Calculation
@@
 This formulation provides a principled way to measure a parameter’s true structural contribution to the final update, moving beyond its simple magnitude.
-Appendix C Description of Baselines
-Our baselines span from foundational FL algorithms to advanced techniques specifically designed for communication-efficient and effective federated fine-tuning of large language models.
-FedAvg (McMahan et al., 2017) is the first algorithm in federated learning.
-It establishes a simple yet effective paradigm for collaborative model training.
-In each communication round, clients download current global model parameters.
-Each client then performs multiple steps of local training on its private dataset.
-The resulting updated model weights are sent back to the server, which aggregates them by computing a weighted average to produce the global model for the subsequent round.
-FedAvg serves as the fundamental baseline for performance and communication cost.
-FedProx (Li et al., 2020) is a key extension to FedAvg designed to address the challenge of statistical heterogeneity (non-IID data) across clients.
-Non-IID data can cause local client models to drift significantly from the global consensus, leading to convergence instability.
-FedProx mitigates this by introducing a proximal term to the local objective function on each client.
-This term regularizes the local training process by penalizing the squared Euclidean distance between the local model weights and the global model weights, thereby constraining local updates and promoting more stable convergence.
-FFA-LoRA (Sun et al., 2024) is a parameter-efficient fine-tuning method tailored for the federated setting, focusing on reducing communication overhead and accurate aggregation.
-It utilizes the Low-Rank Adaptation (LoRA) strategy.
-To achieve communication efficiency, FFA-LoRA freezes the randomly initialized LoRA $A$ matrices across all clients and throughout the training process.
-Consequently, clients only need to train and communicate the $B$ matrices, drastically reducing the number of parameters transmitted in each round.
-FlexLoRA (Bai et al., 2024) proposes a more sophisticated aggregation mechanism for federated LoRA fine-tuning.
-It posits that directly averaging the low-rank matrices from different clients may be suboptimal.
-Instead, clients upload their locally trained LoRA matrices.
-The server first reconstructs the full-rank weight updates for each client ($B_{i}A_{i}$) and averages them.
-Subsequently, the server performs Singular Value Decomposition (SVD) to find a new, optimal low-rank approximation, which is then decomposed into new $B$ and $A$ matrices for client downloading.
-FedICU (Liao et al., [n.
-d.]) is an advanced FL method that decomposes the global model into consensus and divergence components to handle heterogeneity better.
-The core mechanism is a consensus-divergence splitting performed on the server-side during aggregation.
-Consensus components and divergence components of all clients are aggregated separately.
-FEdICU further proposes the selective updating strategy that uploads only more specialized parameters for aggregation.
-Fed-DARE addresses the communication bottleneck in FL by integrating the DARE (Drop and Rescale) technique (Yu et al., 2024).
-In this approach, clients sparsify their local LoRA updates by randomly dropping parameters.
-The remaining non-zero elements are then rescaled to preserve the update’s expected magnitude, resulting in highly sparse matrices that significantly reduce the communication payload.
-We adopt a sparsity ratio of 0.9 as recommended by DARE.
-As shown in Figure 2, this random-based approach outperforms magnitude-based sparsification, while higher sparsity ratios lead to performance degradation.
-This justifies our choice and the exclusion of magnitude-based methods as a baseline.
+B.1. Theoretical Analysis of Pruning Error
+In this section, we provide a theoretical justification for the proposed importance score.
+We demonstrate that our greedy pruning strategy, based on the importance score defined in Equation 6, is mathematically equivalent to minimizing the upper bound of the reconstruction error of the full-rank weight update $\Delta W$.
+B.1.1. Error Bound Derivation
+We analyze the approximation error for the term $\Delta W_{B}=\Delta B_{i}^{t}A_{i}^{t-1}$ (the derivation for $\Delta A$ is symmetric).
+Let $\mathcal{S}$ denote the set of indices in $\Delta B_{i}^{t}$ selected for pruning.
+The pruning error matrix $\mathcal{E}$ is the sum of the rank-1 contributions of the removed elements:
+(20) $\mathcal{E}=\sum_{(u,v)\in\mathcal{S}}(\Delta B_{i}^{t})_{uv}\cdot(\mathbf{e}_{u}\otimes\mathbf{a}_{v}^{T}),$
+where $\mathbf{a}_{v}^{T}$ is the $v$-th row of $A_{i}^{t-1}$. By applying the Triangle Inequality to the Frobenius norm of $\mathcal{E}$, we derive the error upper bound:
+(21) $\|\mathcal{E}\|_{F}$ $=\left\|\sum_{(u,v)\in\mathcal{S}}(\Delta B_{i}^{t})_{uv}(\mathbf{e}_{u}\otimes\mathbf{a}_{v}^{T})\right\|_{F}$
+$\leq\sum_{(u,v)\in\mathcal{S}}\underbrace{|(\Delta B_{i}^{t})_{uv}|\cdot\|\mathbf{a}_{v}\|_{2}}_{I(\Delta B_{i}^{t}[u,v])}.$
+The right-hand side is exactly the sum of the importance scores of the pruned parameters.
+Thus, pruning elements with the lowest scores $I(\cdot)$ strictly minimizes the upper bound of the weight update reconstruction error $\|\Delta W\|_{F}$.
+B.1.2. Comparison with Gradient-based Pruning
+Gradient-based pruning methods are impractical for Federated LoRA Fine-tuning for several reasons:
+  • Computational and Memory Overhead: Calculating the exact influence of parameters on the loss usually requires second-order derivatives (e.g., Hessian).
+For LLMs, computing or even approximating the Hessian is prohibitively expensive for resource-constrained edge devices.
+Even first-order methods require retaining gradients, doubling the memory footprint during the scoring phase.
+  • Federated Constraints: In a federated setting, data is non-IID.
+A gradient-based score computed on local data may not accurately reflect the parameter’s importance to the global model, potentially leading to overfitting local distributions.
+In contrast, our proposed importance score is a data-free and zero-overhead metric.
+It relies solely on the magnitude of the weights and the structural correlations within the LoRA matrices ($\Delta B$ and $A$), which are already present in memory.
+This ensures that our method minimizes the deviation in the weight space ($\|\Delta W\|$) efficiently, offering a superior trade-off between communication efficiency and model performance.
 Table 7. Average In-domain Performance Comparison
 Performance
-FedAvg 33.9
+FedAvg 33.90
    with Random sparsification (i.e., Fed-DARE) 33.57
@@
 Ours (FedSRD / FedSRD-e) $\approx$31.0 $\approx$43.0 $\approx$74.0
-Appendix D Communication Cost Analysis
+Appendix C Communication Cost Analysis
 We analyze the communication cost per round for each method, using a Llama3.2-3B model fine-tuned with LoRA (rank $r=64$) as a case study.
@@
 The overhead from bitmaps can be further mitigated using lossless compression techniques like Run-Length Encoding (Golomb, 1966).
-Appendix E Importance-aware Sparsification Suits Well for FedAvg
+Appendix D Importance-aware Sparsification Suits Well for FedAvg
 To validate the superiority of our proposed importance-aware sparsification over others, we plug and compare them on FedAvg in Table 7.
 The results show that our importance-aware sparsification is more effective in the FL system.
-Appendix F Justification for Equation 7
+Appendix E Justification for Equation 7
 Equation 7 proposes a data-driven approach to dynamically determine the layer-wise sparsity ratio, $\rho$, for the LoRA update matrices $\Delta A_{i}^{t}$ and $\Delta B_{i}^{t}$.
@@
 We leave other possible designs for adaptive determination of sparsity ratio for future work.
-Appendix G Magnitude of $B^{t-1},A^{t-1}$ and $\Delta B^{t},\Delta A^{t}$
+Appendix F Magnitude of $B^{t-1},A^{t-1}$ and $\Delta B^{t},\Delta A^{t}$
 In the Taylor expansion in Equation 11, the product term $\Delta B^{t}\Delta A^{t}$ is treated as a negligible higher-order term.
@@
 This conclusion is further reinforced by the observed convergence trend, where both relative ratios systematically decrease as the number of federated rounds increases.
-Appendix H Performance of Base Models and Locally Trained Models
-We test the performance of base models (Llama3.2-3B and Qwen2-7B) on our benchmarks.
-We also conduct local training on each client (e.g., Code indicates the model is locally trained on the client from the code domain) and test the locally trained models on each client on our benchmarks.
-The results are reported in Table 10.
 Table 9. Performance of base models and local training
@@
 Math 19.25 18.80 43.83 34.16
+Appendix G Performance of Base Models and Locally Trained Models
+We test the performance of base models (Llama3.2-3B and Qwen2-7B) on our benchmarks.
+We also conduct local training on each client (e.g., Code indicates the model is locally trained on the client from the code domain) and test the locally trained models on each client on our benchmarks.
+The results are reported in Table 9.